OpenStudy (anonymous):

16. Choose the ordered pair that represents the solution to the system of equations. 3x - 2y = 10 5x + 3y = 4 (2,2) (-2,2) (-2,-2) (2,-2)

OpenStudy (angela210793):

shumezojme me 3 lart dhe 2 poshte 3x - 2y = 10 (*3)------->9x-6y=30 5x + 3y = 4 (*2) 10x+6y==8(mbledhim ane per ane)-->19x=38 x=38/19 x=2 3*2-2y=10 6-2y=10 -2y=4 y=-2 zgjidhja eshte (2;-2)

OpenStudy (anonymous):

je vertete nje gjeni i rralle angela :) te lumte dora

OpenStudy (anonymous):

system: 3x - 2y = 10 5x + 3y = 4 equal: (2, -2)

OpenStudy (angela210793):

Hahaha :P:P jo s'jam..por ti duhet te lexosh mesimin e matematikes :P:P:P

OpenStudy (anonymous):

jo je he si thua ashtu ?

OpenStudy (anonymous):

"jo je he si thua ashtu ? English please :)"

OpenStudy (anonymous):

sry man, but its our language :)

OpenStudy (angela210793):

best language ever :):) :P

OpenStudy (anonymous):

ok. this system function is very simple, easy.